OpenStudy (anonymous):

Thirty decreased by one half o f a number is one sixth of the square of the number. Find the numbers.

OpenStudy (anonymous):

30 - 1/2(x) = 1/6(x^2)

OpenStudy (lgbasallote):

one half of a number means x/2 one sixth of the square of the number means \(\frac{x^2}{6}\) therefore it's just \[30 - \frac{x}{2} = \frac{x^2}{6}\]

OpenStudy (anonymous):

Solve this equation.

OpenStudy (lgbasallote):

you can multiply the whole equation by the LCM to get rid of the denominators

OpenStudy (anonymous):

Can you find it by using this method?